OpenStudy (anonymous):

3. Simplify the expression. 9xy2 – 11xy2 I combine like terms 9xy^4 rounded to 66xy 66xy/-11 = -6 I have no clue to how to solve this.

OpenStudy (abdullahm):

Is the question: \(\sf\Large 9xy^2-11xy^2=?\)

OpenStudy (anonymous):

Yes

OpenStudy (abdullahm):

we can factor out the \(\sf\Large xy^2\) to make it easier. :)

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

So it's 9 -11 = 3?

OpenStudy (abdullahm):

Use a calculator 9 - 11 = -2

OpenStudy (abdullahm):

It seems like you have slight difficulties in addition/subtraction when the answer is negative...

OpenStudy (anonymous):

Hehe yeah

OpenStudy (anonymous):

so It's xy = -2?

OpenStudy (abdullahm):

one sec

OpenStudy (abdullahm):

\(\sf\large 9xy^2-11xy^2= xy^2(9-11)=xy^2(-2)=-2xy^2\)

OpenStudy (anonymous):

I don't think i'm going to remember that with out going over it a lot of time. Thanks a lot though i appreciate you putting up with me

OpenStudy (abdullahm):

just remember this, it will make it easier \(\sf\Large ab + ac = a(b+c)\)

OpenStudy (anonymous):

Alright, i'll try to keep that in mind.

OpenStudy (anonymous):

Alright, i'll try to keep that in mind.

OpenStudy (abdullahm):

if you forget it, it's called the distributive property. and I just switched it so that it would be helpful for this question. The distributive property actually says: \(\sf\Large a(b+c) = ab+ac\)